# A Learnable Wavelet Transformer for Long-Short Equity Trading and Risk-Adjusted Return Optimization

### v2

@@ Abstract
Learning profitable intraday trading policies from financial time series is challenging due to heavy noise, non-stationarity, and strong cross-sectional dependence among related assets. We propose WaveLSFormer, a learnable wavelet-based long-short Transformer that jointly performs multi-scale decomposition and return-oriented decision learning. Specifically, a learnable wavelet front-end generates low-/high-frequency components via an end-to-end trained filter bank, guided by spectral regularizers that encourage stable and well-separated frequency bands. To fuse multi-scale information, we introduce a low-guided high-frequency injection (LGHI) module that refines low-frequency representations with high-frequency cues while controlling training stability. The model outputs a portfolio of long/short positions that is rescaled to satisfy a fixed risk budget and is optimized directly with a trading objective and risk-aware regularization. Extensive experiments on five years of hourly data across six industry groups, evaluated over ten random seeds, demonstrate that WaveLSFormer consistently outperforms MLP, LSTM and Transformer backbones, with and without fixed discrete wavelet front-ends. On average in all industries, WaveLSFormer achieves a cumulative overall strategy return of $0.607\pm 0.045$ and a Sharpe ratio of $2.157\pm 0.166$, substantially improving both profitability and risk-adjusted returns over the strongest baselines.
I Introduction
@@
Financial economists have long debated whether asset prices fully reflect all available information. The Efficient Market Hypothesis (EMH) suggests that, under rational expectations and frictionless markets, risk-adjusted excess returns should be unpredictable [1]. Nevertheless, decades of empirical evidence indicate that certain technical rules and cross-sectional patterns can yield economically meaningful profits in practice [2]. These observations motivate systematic methods for extracting tradable signals from historical prices and related market features.
@@
With modern computing and large historical datasets, deep learning has become a central tool for modeling complex dependencies in financial time series. A wide range of architectures, from deep feed-forward networks to recurrent models, have been explored for stock prediction and trading [3, 4, 5, 6]. Despite promising results, a key challenge remains: many pipelines are formulated as forecasting problems optimized by point-wise prediction losses, whereas real trading performance is determined by sequential position decisions and their accumulated P&L over time. This objective mismatch can produce models that appear strong under regression/classification metrics yet fail to deliver robust out-of-sample strategy returns, especially when signals are weak and regimes shift.
@@
Financial time series further complicate end-to-end learning. Returns are noisy, heavy-tailed, and heteroskedastic, and exploitable structure may only be visible at certain temporal scales. Moreover, assets within the same market segment are not independent: co-movements and lead-lag effects are prevalent, so effective decisions often require jointly modeling a group of related assets rather than treating each series in isolation [4, 6]. These characteristics motivate models that can extract stable multi-scale representations while aligning training with trading-oriented objectives.
@@
Wavelet transforms are a natural candidate for multi-scale analysis because they provide a time-frequency decomposition that separates slow-moving trends from transient fluctuations. Prior studies often apply fixed discrete wavelet transforms as preprocessing before downstream predictors [7, 8, 9, 10], and wavelet-integrated deep architectures have been surveyed more broadly [11]. In parallel, Transformers have gained popularity for time-series modeling, and wavelet-Transformer hybrids have been proposed to enhance forecasting by combining multi-scale decomposition with attention [12, 13, 14]. However, for trading, two limitations are particularly relevant: many hybrids still rely on fixed wavelets and emphasize forecasting accuracy rather than trading performance, and cross-frequency interactions are often handled implicitly without explicitly controlling how high-frequency information should refine low-frequency representations for stable decision learning.
@@
II Related Work
II-A Forecasting Models and Trading Objectives
Deep learning has been widely adopted for financial time series due to its ability to capture nonlinear dynamics, regime shifts, and complex temporal dependencies. Representative applications include CNN/LSTM-style predictors for stock forecasting [15, 16, 17, 18, 19]. However, a recurring challenge is the objective mismatch between training and deployment. Many methods optimize point-wise forecast losses like MSE or MAE, while trading performance depends on sequential position decisions, constraints, and risk exposure. As a result, sequence predictors do not necessarily intend to achieve robust strategies [20]. Transformers [21] have also become a dominant sequential modeling architecture and have been adapted to time-series forecasting, with influential variants such as Informer [22] and increasingly use in financial prediction [23, 24]. Yet many Transformer-based financial studies still evaluate primarily under forecasting metrics and then apply external trading rules to obtain positions, which can weaken end-to-end alignment with economic utility.
II-B Wavelet Transformer and Frequency Fusion
Financial signals are noisy and non-stationary, motivating wavelet transforms as a time-frequency tool for denoising and multi-scale decomposition before prediction. Classic wavelet neural hybrids and wavelet-reconstruction pipelines, including combinations with ARIMA/LSTM or related deep models, have been explored to improve forecast accuracy on volatile markets [25, 26, 7, 8, 9, 10, 27], and wavelet features have also been used in trading-oriented settings such as reinforcement learning [28]. More recently, wavelet Transformer hybrids construct multi-scale inputs and use attention to aggregate patterns across time and scales [12, 13, 14]. In finance, Stockformer [29] is a recent attempt to integrate Transformer-style modeling with trading-oriented outputs, and surveys summarize wavelet-integrated deep networks for noisy time series [11]. A key modeling question is how information should flow across frequencies: simple concatenation or implicit fusion can allow unstable high-frequency components to dominate learning under noisy financial signals. In this work, we address both the objective mismatch and the frequency-fusion challenge by learning an end-to-end wavelet filter bank jointly with a trading-oriented objective and adopting a low-guided high-frequency injection mechanism, where attention is computed from the low-frequency branch and high-frequency values are injected through a gated residual pathway for stable fusion. Empirically, we benchmark carefully matched MLP/LSTM/Transformer backbones with and without fixed or learnable wavelet front-ends under the same trading protocol.
III Problem Formulation
@@
III-A Input Windows and Model Outputs
@@
$\mathbf{r}_{t}=(r_{1,t},\ldots,r_{d,t})^{\top},\qquad\boldsymbol{\ell}_{t}=(\ell_{1,t},\ldots,\ell_{d,t})^{\top}\in\mathbb{R}^{d}.$
@@
$\mathbf{X}_{t}=\bigl[\boldsymbol{\ell}_{t-L+1},\dots,\boldsymbol{\ell}_{t}\bigr]\in\mathbb{R}^{d\times L},$
@@
Given $\mathbf{X}_{t-1}$, the model outputs a scalar trading signal
$p_{t}=f_{\theta}(\mathbf{X}_{t-1})\in\mathbb{R},$
where $f_{\theta}$ is WaveLSFormer. We map $p_{t}$ to a bounded position via
@@
As discussed in Section III-E, using $w_{t}$ directly can confound signal quality with aggressiveness. We therefore apply a global risk-budget scaling calibrated on the validation set and define the executable position as
$\tilde{w}_{t}=h(w_{t})\in[\tilde{w}_{\min},\tilde{w}_{\max}],$
where $h(\cdot)$ is a fixed piecewise-linear mapping that enforces a prescribed average leverage and clips extremes. In long-short experiments we set $\tilde{w}_{\min}=-L$ and $\tilde{w}_{\max}=L$, while in long-only or short-only settings one bound is set to zero. Section III-E gives the exact form of $h(\cdot)$.
III-B Return Representation
@@ III-C Optimization Target
$\max_{\theta}\;\Bigl({ROI}_{\mathcal{T}}(\theta),\;\mathcal{S}_{\mathcal{T}}(\theta)\Bigr),$
The exact loss components and their implementation details are given in Section VI-A.
III-D Wavelet-based Signal Noise Decomposition
At short horizons, financial returns exhibit low signal-to-noise ratios. Following wavelet denoising, we decompose the target log return into a slowly varying component and a rapidly fluctuating residual,
$\ell_{j^{\star},t}=s_{t}+n_{t},$
where $s_{t}$ captures low-frequency structure (e.g., trend) and $n_{t}$ aggregates high-frequency, largely unpredictable variations. Over an interval $\mathcal{T}$, we define the effective SNR as
@@ III-D Wavelet-based Signal Noise Decomposition
In WaveLSFormer, a neural wavelet front-end performs a learnable time-frequency decomposition of the raw log-return sequence $\{\ell_{j^{\star},t}\}$ into approximations of $\{s_{t}\}$ and $\{n_{t}\}$ using finite-impulse-response low-/high-pass filters. We learn filters that increase $\mathrm{SNR}$ relative to the raw series while producing features that benefit the downstream trading objective. To this end, we add frequency-domain regularizers that promote low-/high-frequency separation and encourage the filter pair to behave approximately as a tight frame; details are given in Section V.
Given multivariate histories $\mathbf{X}_{t-1}$ and the next-step return $\ell_{j^{\star},t}$, our goal is to learn a mapping $f_{\theta}$ that outputs trading signals $p_{t}$ whose induced positions $w_{t}$ maximize the performance of the risk-adjusted strategy, aided by the wavelet front-end that improves the effective SNR of the input. In the following chapters, we will simplify $\ell_{j^{\star},t}$ to $\ell_{t}$.
III-E Position Rule Under a Fixed Risk Budget
Equation (2) bounds the trading signal, so a single trade never exceeds full long/short exposure. Yet different models can induce very different distributions of $w_{t}$: some outputs saturate near $\pm 1$, while others concentrate around 0. Using $w_{t}$ directly would therefore confound signal quality with aggressiveness, since strategies with larger average $|w_{t}|$ effectively take larger bets and thus realize larger gains and losses.
To normalize risk across models, we estimate a global scale on the validation set $\mathcal{D}_{\mathrm{val}}$ with length $T_{\mathrm{val}}=|\mathcal{D}_{\mathrm{val}}|$. We compute the mean absolute magnitude
$s_{\mathrm{val}}=\frac{1}{T_{\mathrm{val}}}\sum_{t\in\mathcal{D}_{\mathrm{val}}}|w_{t}|,$
and rescale the signal using Eq. (8):
$\hat{w}_{t}=\frac{w_{t}}{s_{\mathrm{val}}},$ (8)
so the validation-period average of $|\hat{w}_{t}|$ is close to one.
To suppress tiny, noise-dominated trades, we apply a dead zone with threshold $\tau>0$ and set $\tau=0.01$ in all experiments. We then enforce an optional leverage cap and scale by a target average leverage $L>0$ (we use $L=1$), yielding the executable position in Eq. (9):
@@ III-E Position Rule Under a Fixed Risk Budget
By construction,
@@
up to sparsification from the dead zone, so models are evaluated under a comparable risk budget. The scale $s_{\mathrm{val}}$ is estimated once on $\mathcal{D}_{\mathrm{val}}$ and kept fixed for testing and online simulation, ensuring a consistent and implementable mapping from network outputs to positions at a pre-specified risk level.
IV Data Construction Pipeline
@@ IV-A Data Collection
We obtain sector and industry classifications from StockAnalysis [^1] https://stockanalysis.com/, which publishes curated U.S. stock lists by sector and industry. We extract symbols for each industry and use this stock-industry mapping to define the cross-sectional universe and the industry-group experiments in Section VII.
IV-B Data Preprocessing
@@ IV-C Stock Selection
IV-D Financial Securities with Granger-type Causality
We assume that same-industry stocks exhibit co-movements and lead-lag effects. We quantify such dependencies using nonparametric Granger causality [30], where $x_{t}$ Granger-causes $y_{t}$ if adding lags of $x_{t}$ significantly improves predicting $y_{t}$ beyond using lags of $y_{t}$ alone.
For each industry universe, we run pairwise tests on the log-change series of candidate securities to obtain a directional p-value matrix. We then apply Benjamini-Hochberg FDR control and retain securities whose adjusted p-values with respect to the target asset are below $0.05$ in at least one direction. The retained securities form the final multi-asset input set. Fig. 2 illustrates the selected stocks in Renewable Energy, and Table I reports the full adjusted p-value matrix, which reveals the directed dependence structure within each industry universe.
@@ IV-D Financial Securities with Granger-type Causality
TABLE I: BH-FDR-adjusted $p$-value matrix for pairwise tests on hourly log-percent-change series. Smaller values indicate stronger evidence of dependence. Bold entries denote significant pairs under FDR control ($p_{\text{adj}}<0.05$); bold tickers indicate the retained assets.
[…]

### v3

# arXiv:2601.13435v3[cs.LG] 02 Mar 2026

@@ Abstract
Learning profitable intraday trading policies from financial time series is challenging due to heavy noise, non-stationarity, and strong cross-sectional dependence among related assets. We propose WaveLSFormer, a learnable wavelet-based long-short Transformer that jointly performs multi-scale decomposition and return-oriented decision learning. Unlike standard time-series forecasting that optimizes prediction error and typically requires a separate position-sizing or portfolio-construction step, our model directly outputs a market-neutral long/short portfolio and is trained end-to-end on a trading objective with risk-aware regularization. Specifically, a learnable wavelet front-end generates low-/high-frequency components via an end-to-end trained filter bank, guided by spectral regularizers that encourage stable and well-separated frequency bands. To fuse multi-scale information, we introduce a low-guided high-frequency injection (LGHI) module that refines low-frequency representations with high-frequency cues while controlling training stability. The model outputs a portfolio of long/short positions that is rescaled to satisfy a fixed risk budget and is optimized directly with a trading objective and risk-aware regularization. Extensive experiments on five years of hourly data across six industry groups, evaluated over ten random seeds, demonstrate that WaveLSFormer consistently outperforms MLP, LSTM and Transformer backbones, with and without fixed discrete wavelet front-ends. On average in all industries, WaveLSFormer achieves a cumulative overall strategy return of $0.607\pm 0.045$ and a Sharpe ratio of $2.157\pm 0.166$, substantially improving both profitability and risk-adjusted returns over the strongest baselines.
Keywords: neural wavelet regularization; wavelet-transformer network; low-guided high-frequency injection; return optimization
1 Introduction
@@
Financial economists have long debated whether asset prices fully reflect all available information. The Efficient Market Hypothesis (EMH) suggests that, under rational expectations and frictionless markets, risk-adjusted excess returns should be unpredictable Fama (1970). Nevertheless, decades of empirical evidence indicate that certain technical rules and cross-sectional patterns can yield economically meaningful profits in practice Brock et al. (1992). These observations motivate systematic methods for extracting tradable signals from historical prices and related market features.
@@
With modern computing and large historical datasets, deep learning has become a central tool for modeling complex dependencies in financial time series. A wide range of architectures, from deep feed-forward networks to recurrent models, have been explored for stock prediction and trading Kohzadi et al. (1996); Ozbayoglu et al. (2020); Zhang et al. (2024a); Giantsidi and Tarantola (2025). Despite promising results, a key challenge remains: many pipelines are formulated as forecasting problems optimized by point-wise prediction losses, whereas real trading performance is determined by sequential position decisions and their accumulated P&L over time. This objective mismatch can produce models that appear strong under regression/classification metrics yet fail to deliver robust out-of-sample strategy returns, especially when signals are weak and regimes shift.
@@
Financial time series further complicate end-to-end learning. Returns are noisy, heavy-tailed, and heteroskedastic, and exploitable structure may only be visible at certain temporal scales. Moreover, assets within the same market segment are not independent: co-movements and lead-lag effects are prevalent, so effective decisions often require jointly modeling a group of related assets rather than treating each series in isolation Ozbayoglu et al. (2020); Giantsidi and Tarantola (2025). These characteristics motivate models that can extract stable multi-scale representations while aligning training with trading-oriented objectives.
@@
Wavelet transforms are a natural candidate for multi-scale analysis because they provide a time-frequency decomposition that separates slow-moving trends from transient fluctuations. Prior studies often apply fixed discrete wavelet transforms as preprocessing before downstream predictors Peng et al. (2021a); Rao et al. (2022); Zhang et al. (2024b); Ziółkowski (2025), and wavelet-integrated deep architectures have been surveyed more broadly Wu et al. (2025). In parallel, Transformers have gained popularity for time-series modeling, and wavelet-Transformer hybrids have been proposed to enhance forecasting by combining multi-scale decomposition with attention Sasal et al. (2022); Wei et al. (2025); Ping et al. (2025). However, for trading, two limitations are particularly relevant: many hybrids still rely on fixed wavelets and emphasize forecasting accuracy rather than trading performance, and cross-frequency interactions are often handled implicitly without explicitly controlling how high-frequency information should refine low-frequency representations for stable decision learning.
@@
2 Related Work
2.1 Forecasting Models and Trading Objectives
Deep learning has been widely adopted for financial time series due to its ability to capture nonlinear dynamics, regime shifts, and complex temporal dependencies. Representative applications include CNN/LSTM-style predictors for stock forecasting Hoseinzade and Haratizadeh (2019); Nelson et al. (2017); Moghar and Hamiche (2020); Bao et al. (2017); Buczyński et al. (2023). However, a recurring challenge is the objective mismatch between training and deployment. Many methods optimize point-wise forecast losses like MSE or MAE, while trading performance depends on sequential position decisions, constraints, and risk exposure. As a result, sequence predictors do not necessarily intend to achieve robust strategies Krauss et al. (2017). Transformers Vaswani et al. (2017) have also become a dominant sequential modeling architecture and have been adapted to time-series forecasting, with influential variants such as Informer Zhou et al. (2021) and increasingly use in financial prediction Mozaffari and Zhang (2024); Kabir et al. (2025). Yet many Transformer-based financial studies still evaluate primarily under forecasting metrics and then apply external trading rules to obtain positions, which can weaken end-to-end alignment with economic utility.
2.2 Wavelet Transformer and Frequency Fusion
Financial signals are noisy and non-stationary, motivating wavelet transforms as a time-frequency tool for denoising and multi-scale decomposition before prediction. Classic wavelet neural hybrids and wavelet-reconstruction pipelines, including combinations with ARIMA/LSTM or related deep models, have been explored to improve forecast accuracy on volatile markets Lei (2018); Peng et al. (2021b, a); Rao et al. (2022); Zhang et al. (2024b); Ziółkowski (2025); Chandar et al. (2016), and wavelet features have also been used in trading-oriented settings such as reinforcement learning Lee et al. (2021). More recently, wavelet Transformer hybrids construct multi-scale inputs and use attention to aggregate patterns across time and scales Sasal et al. (2022); Wei et al. (2025); Ping et al. (2025). In finance, Stockformer Ma et al. (2025) is a recent attempt to integrate Transformer-style modeling with trading-oriented outputs, and surveys summarize wavelet-integrated deep networks for noisy time series Wu et al. (2025). A key modeling question is how information should flow across frequencies: simple concatenation or implicit fusion can allow unstable high-frequency components to dominate learning under noisy financial signals. In this work, we address both the objective mismatch and the frequency-fusion challenge by learning an end-to-end wavelet filter bank jointly with a trading-oriented objective and adopting a low-guided high-frequency injection mechanism, where attention is computed from the low-frequency branch and high-frequency values are injected through a gated residual pathway for stable fusion. Empirically, we benchmark carefully matched MLP/LSTM/Transformer backbones with and without fixed or learnable wavelet front-ends under the same trading protocol.
3 Problem Formulation
@@
3.1 Input Windows and Model Outputs
@@
$\mathbf{r}_{t}=(r_{1,t},\ldots,r_{d,t})^{\top},\qquad\bm{\ell}_{t}=(\ell_{1,t},\ldots,\ell_{d,t})^{\top}\in\mathbb{R}^{d}.$
@@
$\mathbf{X}_{t}=\bigl[\bm{\ell}_{t-L+1},\dots,\bm{\ell}_{t}\bigr]\in\mathbb{R}^{d\times L},$
@@
Given $\mathbf{X}_{t-1}$, the model outputs a scalar trading signal $p_{t}=f_{\theta}(\mathbf{X}_{t-1})\in\mathbb{R}$, where $f_{\theta}$ is WaveLSFormer. We map $p_{t}$ to a bounded position via
@@
As discussed in Section 3.5, using $w_{t}$ directly can confound signal quality with aggressiveness. We therefore apply a global risk-budget scaling calibrated on the validation set and define the executable position as $\tilde{w}_{t}=h(w_{t})\in[\tilde{w}_{\min},\tilde{w}_{\max}]$, where $h(\cdot)$ is a fixed piecewise-linear mapping that enforces a prescribed average leverage and clips extremes. In long-short experiments we set $\tilde{w}_{\min}=-L$ and $\tilde{w}_{\max}=L$, while in long-only or short-only settings one bound is set to zero. Section 3.5 gives the exact form of $h(\cdot)$.
3.2 Return Representation
@@ 3.3 Optimization Target
$\theta^{*}=\arg\max_{\theta}\;\Bigl({ROI}_{\mathcal{T}}(\theta),\;\mathcal{S}_{\mathcal{T}}(\theta)\Bigr),$
The exact loss components and their implementation details are given in Section 6.1.
3.4 Wavelet-based Signal Noise Decomposition
At short horizons, financial returns exhibit low signal-to-noise ratios. Following wavelet denoising, we decompose the target log return into a slowly varying component and a rapidly fluctuating residual $\ell_{j^{\star},t}=s_{t}+n_{t}$, where $s_{t}$ captures low-frequency structure and $n_{t}$ aggregates high-frequency, largely unpredictable variations. Over an interval $\mathcal{T}$, we define the effective SNR as
@@ 3.4 Wavelet-based Signal Noise Decomposition
In WaveLSFormer, a neural wavelet front-end performs a learnable time-frequency decomposition of the raw log-return sequence $\{\ell_{j^{\star},t}\}$ into approximations of $\{s_{t}\}$ and $\{n_{t}\}$ using finite-impulse-response low-/high-pass filters. We learn filters that increase $\mathrm{SNR}$ relative to the raw series while producing features that benefit the downstream trading objective. To this end, we add frequency-domain regularizers that promote low-/high-frequency separation and encourage the filter pair to behave approximately as a tight frame; details are given in Section 5.
Given multivariate histories $\mathbf{X}_{t-1}$ and the next-step return $\ell_{j^{\star},t}$, our goal is to learn a mapping $f_{\theta}$ that outputs trading signals $p_{t}$ whose induced positions $w_{t}$ maximize the performance of the risk-adjusted strategy, aided by the wavelet front-end that improves the effective SNR of the input. In the following chapters, $\ell_{j^{\star},t}$ is simplified to $\ell_{t}$.
3.5 Position Rule Under a Fixed Risk Budget
Eq. (2) bounds the trading signal, so a single trade never exceeds full long/short exposure. Yet different models can induce very different distributions of $w_{t}$. Some outputs saturate near $\pm 1$, while others concentrate around 0. Using $w_{t}$ directly would therefore confound signal quality with aggressiveness, since strategies with larger average $|w_{t}|$ effectively take larger bets and thus realize larger gains and losses.
To normalize risk across models, we estimate a global scale on the validation set $\mathcal{D}_{\mathrm{val}}$ with length $T_{\mathrm{val}}=|\mathcal{D}_{\mathrm{val}}|$. We compute the mean absolute magnitude and rescale the signal using $\hat{w}_{t}$, so the validation-period average of $|\hat{w}_{t}|$ is close to one.
$s_{\mathrm{val}}=\frac{1}{T_{\mathrm{val}}}\sum_{t\in\mathcal{D}_{\mathrm{val}}}|w_{t}|,\qquad\hat{w}_{t}=\frac{w_{t}}{s_{\mathrm{val}}},$ (8)
To suppress tiny, noise-dominated trades, we apply a dead zone with threshold $\tau>0$ and set $\tau=0.01$ in all experiments. We then enforce an optional leverage cap and scale by a target average leverage $L>0$ (we use $L=1$), yielding the executable position in
@@ 3.5 Position Rule Under a Fixed Risk Budget
By this construction, the position is up to sparsification from the dead zone, so models are evaluated under a comparable risk budget. The scale $s_{\mathrm{val}}$ is estimated once on $\mathcal{D}_{\mathrm{val}}$ and kept fixed for testing and online simulation, ensuring a consistent and implementable mapping from network outputs to positions at a pre-specified risk level.
@@
4 Data Construction Pipeline
@@ 4.1 Data Collection
We obtain sector and industry classifications from StockAnalysis [^1] https://stockanalysis.com/, which publishes curated U.S. stock lists by sector and industry. We extract symbols for each industry and use this stock-industry mapping to define the cross-sectional universe and the industry-group experiments in Section 7.
4.2 Data Preprocessing
@@ 4.3 Stock Selection
4.4 Financial Securities with Granger-type Causality
We assume that same-industry stocks exhibit co-movements and lead-lag effects. We quantify such dependencies using nonparametric Granger causality Diks and Panchenko (2006), where $x_{t}$ Granger-causes $y_{t}$ if adding lags of $x_{t}$ significantly improves predicting $y_{t}$ beyond using lags of $y_{t}$ alone.
For each industry universe, we run pairwise tests on the log-change series of candidate securities to obtain a directional p-value matrix. We then apply Benjamini-Hochberg FDR control and retain securities whose adjusted p-values with respect to the target asset are below $0.05$ in at least one direction. The retained securities form the final multi-asset input set. Fig. 2 illustrates the selected stocks in Renewable Energy, and Table 1 reports the full adjusted p-value matrix, which reveals the directed dependence structure within each industry universe.
@@ 4.4 Financial Securities with Granger-type Causality
Table 1: BH-FDR-adjusted $p$-value matrix for pairwise tests on hourly log-percent-change series. Smaller values indicate stronger evidence of dependence. Bold entries denote significant pairs under FDR control ($p_{\text{adj}}<0.05$); bold tickers indicate the retained assets.
AEP ATO AWK CMS CNP CWEN EIX EVRG IDA NEE OTTR PEG PNW POR WEC XEL